OpenStudy (anonymous):

the set of values of a for which \[x^2+ax+\sin^{-1} (x^2-4x+5)+\cos^{-1} (x^2-4x+5)\]has at least one real solution is

OpenStudy (anonymous):

@geerky42 ? @annas ? @beccamarx19 ? @Conqueror ? @Destinymasha ? @elementwielder ? @Future_UsArmy_MP ? @ganeshie8 ?

OpenStudy (anonymous):

i think this would be simplified to

OpenStudy (anonymous):

i think this would be simplified to \[x^2+ax+\pi/2=0\]

OpenStudy (anonymous):

now how to solve further?

OpenStudy (anonymous):

should i apply b^2-4ac>=0

OpenStudy (anonymous):

bu the answer is not coming correct if i apply it!!!

OpenStudy (anonymous):

@mathmale ?

OpenStudy (anonymous):

@naylah ? @blues ?

OpenStudy (anonymous):

What you have given is an expression. I cannot seen why an expression has a solution.

OpenStudy (anonymous):

Should there be another condition that \[x ^{2}-4x+5\le1 \] ie x=2. Thus a=-8-pi/4